OpenStudy (anonymous):

what is the nth term in this pattern? term: 1 2 3 4 5 6 Value: 4 3 1 -1 -3 -5

OpenStudy (amistre64):

notice that each term is simply the one before it - 2

OpenStudy (amistre64):

... that was almost right llll

OpenStudy (amistre64):

you sure you dint typo the first value?

OpenStudy (amistre64):

A{2} = 3 - 2(0) A{3} = 3 - 2(1) A{4} = 3 - 2(2) A{5} = 3 - 2(3) .............. A{n} = 3 - 2(n-2) with the restriction that A{1} = 4